Suppose three coins are tossed. Let S be the random variable representing





the number of heads that occur. Determine the sample space and the values of a





random variable

1
Expert's answer
2022-02-25T11:54:59-0500

Sample space: {HHH, HHT, HTH, THH, HTT, THT, TTH, TTT}

"P(S=0)=\frac{1}{8}."

"P(S=1)=\frac{3}{8}."

"P(S=2)=\frac{3}{8}."

"P(S=3)=\frac{1}{8}."